# Representation Bending for Large Language Model Safety

### v1

@@ 1 Introduction
Large language models (LLMs) have become versatile tools with notable capabilities, showing promise as general-purpose task solvers. Their growing adoption across many industries and for personal use makes it increasingly important to ensure they are safe and do not cause harmful or catastrophic outcomes Hendrycks et al. (2023). LLMs are typically fine-tuned for both instruction-following and safety, to become both helpful and harmless – to obey and provide helpful responses to benign requests, and to refuse harmful ones Bai et al. (2022). Nevertheless, they still exhibit harmful behaviors, especially to adversarial manipulations Wei et al. (2024); Zou et al. (2023b); Andriushchenko and Flammarion (2024); Pelrine et al. (2023), or even fine-tuning Lermen et al. (2023); Zhan et al. (2024); Anwar et al. (2024); Qi et al. (2024). These attacks can bypass safety training of LLMs and result in models that generate harmful responses.
@@
Existing safety training methods (e.g., SFT Brown et al. (2020), DPO Rafailov et al. (2024a), RLHF Christiano et al. (2017)) are frequently bypassed Zou et al. (2023b); Wei et al. (2024); Andriushchenko et al. (2024); Schwinn et al. (2024). To combat these, adversarial training is proposed, where during training specific attack methods are remedied Mazeika et al. (2024); Jiang et al. (2024b). However, adversarial training only addresses specific types of attacks and does not generalize to attacks that were unseen during training. System-level defenses, such as input and output filters Han et al. (2024), may be hard to scale due to the frequency of checks for every query, need to be upgraded with new and intelligent attacks, and do not make the underlying model inherently safer Zou et al. (2024), as they only limit the model’s exposure to unsafe content.
RepBend drives inspiration from these works, and also from the activation steering literature (Turner et al., 2024; Panickssery et al., 2023; Cao et al., 2024). Activation steering forms “steering vectors” by simply taking the difference of the activations for safe and unsafe prompts, and performs simple arithmetic (e.g., addition or subtraction) to change the model’s behavior during inference. Activation steering has limitations for in- and out-of-distribution generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024).
@@ 2 Related Work
Conventional unlearning aims to update the weights of a model to remove specific knowledge, and usually is about narrow topics (e.g., Harry Potter) or fictional information Eldan and Russinovich (2023); Maini et al. (2024). For reviews of existing works, see Liu et al. (2024). Although fine-grained control is useful, our work tackles broader notions of undesirable outputs. In the context of language models, previous work explores concepts like fairness, privacy, safety or hallucinations Jang et al. (2023); Yao et al. (2023); Liu et al. (2024). Recently NPO Zhang et al. (2024) is proposed; a simple alignment-inspired method that could be extended to unlearn hazardous knowledge. In this work we compare NPO’s performance with RepBend.
@@
A general tactic steers language models away from generating undesirable text during inference. Liu et al. (2023) proposes in-context vectors that encode and replace in-context examples and uses it with simple vector arithmetic during inference, while other works control LLMs by “steering vector” activations Turner et al. (2024); Panickssery et al. (2023); Cao et al. (2024). Similarly, Jorgensen et al. (2023) applies mean clustering to find better steering vectors than normal averaging, and Qiu et al. (2024) proposes spectral editing of activations, an activation editing method to guide LLMs to generate desirable outputs through spectral decomposition. Critically, these methods focus on inference-time changes to the forward pass. While this makes activation steering widely applicable, activation steering has limitations for in- and out-of-distribution (OOD) generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024). RepBend has OOD generalizability and good general reasoning score (Section 4.2).
@@ 3 Representation Bending
To do so, we have to bring the model into its unsafe and safe representation states: we input safe and unsafe text to the model, to elicit the targeted representations, and read the activations of the model for those texts Liu et al. (2023); Turner et al. (2024); Zou et al. (2024, 2023a). We first need to gather a dataset consisting of safe and unsafe texts. We need text that is “unsafe and also text that is considered “safe to elicit model’s representations. Details of the datasets we used discussed in Section A.1.
@@
RepBend is shown in Algorithm 1. RepBend tries to change the representations of model by minimizing a loss function over a fixed number of steps during fine-tuning. At a high level, RepBend bends the representations of model to be far from unsafe representations and close to safe representations, while maintaining its general capabilities. For improved performance, RepBend updates the model parameters with LoRA Hu et al. (2021). Model with LoRA parameters is referred to as (and can be seen as an independent) model . Since we only update the LoRA parameters, at the end of the algorithm would be the desired safe model. We now explain all the steps in Algorithm 1.
@@
We sample a batch of safe text from safe sets, , safe prompts, and , unsafe prompts followed by safe responses, respectively, and feed them into the LLM to elicit safe representations (line 5). We then obtain the vector , the difference between safe representations of and (line 6). For simpler notation, we show as , denoting representations of model for a set of layers and a set of token positions . (Prior work typically uses for positions of last tokens of prompt, or first tokens of response, or all tokens of input von Rütte et al. (2024); Gurnee and Tegmark (2023); Zou et al. (2023a); Turner et al. (2024); Panickssery et al. (2023); Zou et al. (2024)). We sample a batch of unsafe text from unsafe set, , and get the unsafe representation difference vector (line 7-8). Since we want the model not to generate unsafe text, we like to have big L2 norm ( be far from for unsafe representation), while we want to have small L2 norm ( not get far from for safe representation). Hence the L2 norm of these vectors are added to the loss term (line 12) with respective negative and positive signs.
@@
Since we also want to have stability while bending the representations, we add two loss terms with distinct purposes: KL divergence loss term between output logits of and is added so that the model still retains its general capability when input is safe. A similarity term denoted by is added to encourage similarity between outputs of unsafe prompts, given the fact that the response of a safe LLM to most unsafe prompts would begin by refusal texts (e.g.,“I’m Sorry, I am unable to assist”, or “As an assistant, I am unable to provide help.”) Arditi et al. (2024); Liu et al. (2023); von Rütte et al. (2024). In lines 9-10, a batch of unsafe text are sampled and are added to a set . Method computes average of cosine similarities between all pairs in the set . Since the batch size is often not big for LLMs, the complexity of this loss is manageable. By minimizing this loss for model , this algorithm makes the LLM safe.
@@
Finding the set of layers to do the intervention of RepBend is non-trivial and largely dependent on the architecture and practical considerations Wu et al. (2024); Arditi et al. (2024); Li et al. (2024b). Prior work has considered early layers Turner et al. (2024); Li et al. (2024a), middle layers Li et al. (2024b); Zou et al. (2024), later layers Zou et al. (2023a); Cao et al. (2024); Panickssery et al. (2023), and all layers Liu et al. (2023) for intervening on activations. We need to find the layers for which representation bending can have the best impact. We hypothesized that mid to later layers are responsible for the generation of output, and targeting them could be good for representation bending. This aligns well with prior work Zou et al. (2023a); Panickssery et al. (2023) that shows behavior clustering and emotion representations emerge around half or one-third of the way through the layers, indicating rich representations. We validated our hypothesis in our experiments too (Section 4.6), where representation bending works best in mid to later layers (layers 20 and after) among other choices.
@@
Following existing research, we apply representation bending in the residual stream of LLM Arditi et al. (2024); Elhage et al. (2021); von Rütte et al. (2024); Yu et al. (2024); Panickssery et al. (2023); Zou et al. (2024), in particular the residual stream at the output of the block . We also experimented with other possibilities when activations are obtained from different parts of the transformer block, namely activations at the input of (after first layer normalization ), after passing through the self-attention block , and at the input of after the second layer normalization. We observed not much variability, and decided to go with residual streams as in prior works.
@@ Results.
Additional results with more algorithms like SFT, DPO Rafailov et al. (2024b), WHP Eldan and Russinovich (2023), and the original model with a safe instructed prompt (e.g., “If someone asks you an unsafe or harmful prompt, do not answer”) are available in Section B.1. Moreover, further comparison with controllable text generation methods (e.g., a classifier-based method Yang and Klein (2021)) with better transferability are discussed in Section B.2
@@
Table 3 shows the performance of RepBend across the additional LLM architectures. With minimal hyperparameter tuning, RepBend successfully enhances safety while maintaining usability and capability. These results confirm the scalability of RepBend across diverse architectures and parameter sizes.
4.5 Impact of cos_sim loss
Results.
@@ 4.6 Internal Behavior of Model
Representation Analysis.
To analyze the impact of RepBend on internal representations, we need a way to visualized the model’s internals. We use the Logit Lens framework Nostalgebraist (2020), which maps model’s representations from the latent space to the vocabulary space for interpretable analysis. Logit Lense hence allows us to see what the next predicted token in each layer is and it is widely used to investigate Transformer-based architectures Dar et al. (2023); Elhage et al. (2021).
@@ 6 Limitations
An inherent limitation of any unlearning method, is robustness to re-learning hazardous knowledge. Similar to the finding for RMU Li et al. (2024a), we also found out that if RepBend is fine-tuned with unsafe data, it can relearn the harmful content. This limitation of robustness has also been observed by several others Qi et al. (2024); Lo et al. (2024); Lermen et al. (2023); Zhan et al. (2024), and is still an ongoing problem in the community Lynch et al. (2024), with some recent work to remedy this Rosati et al. (2024); Lyu et al. (2024). Future work can borrow techniques to make RepBend safe post-tuning.
@@ A.5 Evaluation Benchmarks
Figure 4: Ablation study of cos_sim loss term in RepBend on Llama3 8B Instruct. All other hyperparameters are fixed.
Figure 5: Impact of the loss-weighting hyperparameter on safety, measured as the average ASR on WildGuardTest, HarmBench, and PAP. Each color represent a configuration of hyper-parameters. The cos_sim is relatively unaffected by , with similar trends across different hyperparameter combinations. -
@@
GCG Zou et al. (2023b): updates input tokens to improve possibility of jailbreaking via gradient optimization: by maximizing compliance likelihood, GCG finds an adversarial suffix for a prompt.
@@
Input Embed Schwinn et al. (2024): Produces adversarial input embeddings instead of tokens using gradient-based optimization.
We use 320 HarmBench samples for GCG, while Prefilling and Input Embed are tested on 240 harmful samples following their original implementation We run GCG optimization for 500 iterations, with adversarial string initiated as “! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! !” (default settings of HarmBench), and searching 512 adversarial candidates at each step. The evaluations (except TrustLLM) use the Open-source classifier from HarmBench Mazeika et al. (2024) to determine attack success rates, compliance, and general quality of the responses. TurstLLM-Jailbreak is evaluated with GPT-4.

### v2

@@ 1 Introduction
Large language models (LLMs) have become versatile tools with notable capabilities, showing promise as general-purpose task solvers. Their growing adoption across many industries and for personal use makes it increasingly important to ensure they are safe and do not cause harmful or catastrophic outcomes Hendrycks et al. (2023); Phuong et al. (2024). LLMs are typically fine-tuned for both instruction-following and safety, to become both helpful and harmless – to obey and provide helpful responses to benign requests, and to refuse harmful ones Bai et al. (2022). Nevertheless, they still exhibit harmful behaviors, especially to adversarial manipulations Wei et al. (2024a); Zou et al. (2023b); Andriushchenko and Flammarion (2024); Pelrine et al. (2023), or even fine-tuning Lermen et al. (2023); Zhan et al. (2024); Anwar et al. (2024); Qi et al. (2024b). These attacks can bypass safety training of LLMs and result in models that generate harmful responses.
@@
Existing safety training methods (e.g., SFT Brown et al. (2020), DPO Rafailov et al. (2024a), RLHF Christiano et al. (2017)) are frequently bypassed Zou et al. (2023b); Wei et al. (2024a); Andriushchenko et al. (2024); Schwinn et al. (2024) and are subject to “shallow safety alignment” Qi et al. (2024a). To combat these, adversarial training is proposed, where during training specific attack methods are remedied Mazeika et al. (2024); Jiang et al. (2024b). However, adversarial training only addresses specific types of attacks and does not generalize to attacks that were unseen during training. System-level defenses, such as input and output filters Han et al. (2024), may be hard to scale due to the frequency of checks for every query, need to be upgraded with new and intelligent attacks, and do not make the underlying model inherently safer Zou et al. (2024), as they only limit the model’s exposure to unsafe content.
RepBend drives inspiration from these works, and also from the activation steering literature (Turner et al., 2024; Panickssery et al., 2023; Cao et al., 2024b). Activation steering forms “steering vectors” by simply taking the difference of the activations for safe and unsafe prompts, and performs simple arithmetic (e.g., addition or subtraction) to change the model’s behavior during inference. Activation steering has limitations for in- and out-of-distribution generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024).
@@ 2 Related Work
Recent studies have explored the limitations of alignment techniques in LLMs. Lin et al. (2023) demonstrates that alignment can be bypassed through carefully designed in-context prompts, while Wei et al. (2024b) reveals that even minimal structural changes can significantly degrade alignment, raising concerns about the robustness of safety measures. Hsu et al. (2024) introduces a low-rank adaptation technique that reduces safety risks during model fine-tuning. Further probing latent representations, Ball et al. (2024) analyzes how jailbreaks succeed by exploiting latent space dynamics, while Treutlein et al. (2024) shows that LLMs infer latent information from evidence distributed across training documents, underscoring the complexity of internal model representations and the difficulty of safety monitoring.
@@
Conventional unlearning aims to update the weights of a model to remove specific knowledge, and usually is about narrow topics (e.g., Harry Potter) or fictional information Eldan and Russinovich (2023); Maini et al. (2024); Wang et al. (2024). For reviews of existing works, see Liu et al. (2024). Authors in Maini et al. (2024); Jin et al. (2024); Hong et al. (2024); Shi et al. (2024) propose real-world LLM unlearning benchmarks that consider task settings, knowledge sources, privacy, scalability, and even internal model parameters. Although fine-grained control is useful, our work tackles broader notions of undesirable outputs. In the context of language models, previous work explores concepts like fairness, privacy, safety or hallucinations Jang et al. (2023); Yao et al. (2023); Liu et al. (2024). Recently NPO Zhang et al. (2024) is proposed; a simple alignment-inspired method that could be extended to unlearn hazardous knowledge. In this work we compare NPO’s performance with RepBend.
@@
A general tactic steers language models away from generating undesirable text during inference. Liu et al. (2023) proposes in-context vectors that encode and replace in-context examples and uses it with simple vector arithmetic during inference, while other works control LLMs by “steering vector” activations Turner et al. (2024); Panickssery et al. (2023); Cao et al. (2024b, a). Similarly, Jorgensen et al. (2023) applies mean clustering to find better steering vectors than normal averaging, and Qiu et al. (2024) proposes spectral editing of activations, an activation editing method to guide LLMs to generate desirable outputs through spectral decomposition. Critically, these methods focus on inference-time changes to the forward pass. While this makes activation steering widely applicable, activation steering has limitations for in- and out-of-distribution (OOD) generalizability, and can compromise model’s general reasoning capabilities Tan et al. (2024); Anthropic (Oct 25, 2024). RepBend has OOD generalizability and good general reasoning score (Section 4.2).
@@ 3 Representation Bending
To do so, we have to bring the model into its unsafe and safe representation states: we input safe and unsafe text to the model, to elicit the targeted representations, and read the activations of the model for those texts Liu et al. (2023); Turner et al. (2024); Zou et al. (2024, 2023a). We first need to gather a dataset consisting of safe and unsafe texts. We need text that is “unsafe” and also text that is considered “safe” to elicit model’s representations. Details of the datasets we used discussed in Section A.1.
@@
RepBend is shown in Algorithm 1. RepBend tries to change the representations of model by minimizing a loss function over a fixed number of steps during fine-tuning. At a high level, RepBend bends the representations of model to be far from unsafe representations and close to safe representations, while maintaining its general capabilities. For improved performance, RepBend updates the model parameters with LoRA Hu et al. (2021). Model with LoRA parameters is referred to as (and can be seen as an independent) model . Since we only update the LoRA parameters, at the end of the algorithm would be the desired safe model.
Loss Terms.
Line 12 shows four terms. in the loss function. 1. First loss term keeps the model’s representation close to the safe representations (small L2 difference); this can be seen as “retain loss.” 2. Second loss term is the opposite of the first one: it pushes the representations of the model to be far away from unsafe representations (large L2 difference); this can be seen as “forget loss” somehow. 3. The cosine similarity loss term stabilizes the model responses under the unsafe queries toward a single “refusal-like” direction (e.g., by terms like “I am sorry I cannot help” or “As an AI assistant, I am unable to …”), helping the model consistently reject them rather than producing random or inconsistent outputs. Recall that cosine similarity is high when inputs are similar, and we want similarity, hence the negative sign for the loss. 4. Finally, the KL divergence loss term keeps the model’s general capabilities intact by encouraging its safe outputs to remain aligned with the original distribution.
@@
We now explain all the steps in Algorithm 1. We sample a batch of safe text from safe sets, , safe prompts, and , unsafe prompts followed by safe responses, respectively, and feed them into the LLM to elicit safe representations (line 5). We then obtain the vector , the difference between safe representations of and (line 6). For simpler notation, we show as , denoting representations of model for a set of layers and a set of token positions . (Prior work typically uses for positions of last tokens of prompt, or first tokens of response, or all tokens of input von Rütte et al. (2024); Gurnee and Tegmark (2023); Zou et al. (2023a); Turner et al. (2024); Panickssery et al. (2023); Zou et al. (2024)). We sample a batch of unsafe text from unsafe set, , and get the unsafe representation difference vector (line 7-8). Since we want the model not to generate unsafe text, we like to have big L2 norm ( be far from for unsafe representation), while we want to have small L2 norm ( not get far from for safe representation). Hence the L2 norm of these vectors are added to the loss term (line 12) with respective negative and positive signs.
@@
Since we also want to have stability while bending the representations, we add two loss terms with distinct purposes: KL divergence loss term between output logits of and is added so that the model still retains its general capability when input is safe. A similarity term denoted by is added to encourage similarity between outputs of unsafe prompts, given the fact that the response of a safe LLM to most unsafe prompts would begin by refusal texts (e.g.,“I’m Sorry, I am unable to assist”) Arditi et al. (2024b); Liu et al. (2023); von Rütte et al. (2024). In lines 9-10, a batch of unsafe text are sampled and are added to a set . Method computes average of cosine similarities between all pairs in the set . Since the batch size is often not big for LLMs, the complexity of this loss is manageable. By minimizing this loss for model , this algorithm makes the LLM safe.
@@
Finding the set of layers to do the intervention of RepBend is non-trivial and largely dependent on the architecture and practical considerations Wu et al. (2024); Arditi et al. (2024b); Li et al. (2024b). Prior work has considered early layers Turner et al. (2024); Li et al. (2024a), middle layers Li et al. (2024b); Zou et al. (2024), later layers Zou et al. (2023a); Cao et al. (2024b); Panickssery et al. (2023), and all layers Liu et al. (2023) for intervening on activations. We need to find the layers for which representation bending can have the best impact. We hypothesized that mid to later layers are responsible for the generation of output, and targeting them could be good for representation bending. This aligns well with prior work Zou et al. (2023a); Panickssery et al. (2023) that shows behavior clustering and emotion representations emerge around half or one-third of the way through the layers, indicating rich representations. We validated our hypotheses in our experiments too (Section 4.5 and Section B.4), where representation bending works best in mid to later layers (layers 20 and after).
@@
Following existing research, we apply representation bending in the residual stream of LLM Arditi et al. (2024b); Elhage et al. (2021); von Rütte et al. (2024); Yu et al. (2024); Panickssery et al. (2023); Zou et al. (2024), in particular the residual stream at the output of the block . We also experimented with other possibilities when activations are obtained from different parts of the transformer block, namely activations at the input of (after first layer normalization ), after passing through the self-attention block , and at the input of after the second layer normalization. We observed not much variability, and decided to go with residual streams as in prior works.
@@ Results.
Additional results with more algorithms like SFT, DPO, WHP Eldan and Russinovich (2023), and the original model with a safe instructed prompt (e.g., “If someone asks you an unsafe or harmful prompt, do not answer”) are available in Section B.1. Moreover, further comparison with controllable text generation methods (e.g., a classifier-based method Yang and Klein (2021)) with better transferability are in Section B.5
@@
Table 3 shows the performance of RepBend across the additional LLM architectures. With minimal hyperparameter tuning, RepBend successfully enhances safety while maintaining usability and capability. These results confirm the scalability of RepBend across diverse architectures and parameter sizes. Experiments on more white-box jailbreak attacks, SCAV (Xu et al., 2024), Weight Orthogonalization (Arditi et al., 2024a), and GuidedBench (Huang et al., 2025) are reported in Section B.2.
@@ 4.5 Internal Behavior of Model
To analyze the impact of RepBend on internal representations, we need a way to visualized the model’s internals. We use the Logit Lens framework Nostalgebraist (2020); Belrose et al. (2023), which maps model’s representations from the latent space to the vocabulary space for interpretable analysis. Logit Lense hence allows us to see what the next predicted token in each layer is and it is widely used to investigate Transformer-based architectures Dar et al. (2023); Elhage et al. (2021).
@@
activations of Llama3 8B
Instruct (original weight).
activations of Llama3 8B
Instruct (RepBend).
distance between activations
of safe and unsafe prompts.
@@
4.6 Analysis of Activations for Safe and Unsafe Prompts
Figure 4 shows PCA plots for activations of safe and unsafe prompts at last layer of Llama3 8B Instruct model, and measured mean Euclidean distance and Jensen Shannon divergence between the two activations of safe and unsafe prompts. We randomly chose 1,000 harmful prompts and 1,000 harmless prompts from wildjailbreak Jiang et al. (2024b), extracted activations of those prompts from each hidden layer, and analyzed the obtained activations.
Figure 4(a) shows that in the original Llama model, representations of those harmless and harmful prompts are clustered into safe and unsafe regions in the space. Figure 4(b) illustrates that RepBend bends the representations, breaking the clustered harmless and harmful regions, potentially (but arguably) getting rid of unsafe region.
Since PCA reduces and shows a high dimensional space in a two dimensional graph, we need other visualizations to to obtain more insights. We measured the distance between the distributions of activations of harmful and harmless prompts for RepBend and the original model. Figure 4(c) shows the layer-wise Euclidean distance between the activations of harmful and harmless prompts, and Figure 4(d) illustrates layer-wise Jensen Shannon divergence between the two sets of activations. These two plots indicate that RepBend “pushes apart” representations of unsafe prompts from safe prompts as a result of bent representation space.
In the original model, we can see that the representations of unsafe and safe prompts are near each other, which could mean the model cannot distinguish them, and it (potentially) can answer unsafe, even when the model is instructed to be safe. Conversely, when the model finetuned with RepBend, we can see that the distance between the representations of safe and unsafe inputs is larger, specially at the later layers, suggesting that the model (arguably) can distinguish safe and unsafe prompts. While RepBend distinctly separates harmful from harmless behavior, which increases the likelihood to correctly refuse the harmful requests, it may potentially cause gibberish outputs in some cases, as shown in Section B.8.
@@ 6 Limitations
An inherent limitation of any unlearning method, is robustness to re-learning hazardous knowledge. Similar to the finding for RMU Li et al. (2024a); Łucki et al. (2024), we also found out that if RepBend is fine-tuned with unsafe data, it can relearn the harmful content. This limitation of robustness has also been observed by several others Qi et al. (2024b); Lo et al. (2024); Lermen et al. (2023); Zhan et al. (2024), and is still an ongoing problem in the community Lynch et al. (2024); Shumailov et al. (2024), with some recent work to remedy this Rosati et al. (2024); Lyu et al. (2024). Future work can borrow techniques to make RepBend safe post-tuning.
@@ Acknowledgments
This work was partially funded by an unrestricted gift from Google. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00354218), the Institute of Information & communications Technology Promotion (IITP) grant funded by the Korea government (MSIT) (No. RS-2024-00457882 Artificial Intelligence Research Hub Project), and the Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2025-02263598, Development of Self-Evolving Embodied AGI Platform Technology through Real-World Experience). This work was partly supported by IITP grants (No.RS-2022-II220077, No.RS-2022-II220113, No.RS-2022-II220959, No.RS-2022-II220871, No.RS-2022-II220951 (50%), No.RS-2021-II211343 (SNU AI), No.RS-2021-II212068 (AI Innovation Hub)) funded by the Korea government(MSIT) and the BK21 FOUR program, SNU in 2025.
@@ A.5 Evaluation Benchmarks
GuidedBench Huang et al. (2025): A benchmark designed to evaluate jailbreak attacks by assessing model responses under guided decoding. It includes 3,000 adversarial prompts spanning 30 harmful intent categories. It uses a set of automatic and human evaluation metrics such as harmfulness, helpfulness, and intent alignment to measure how successfully the model’s outputs comply with adversarial instructions. Outputs are judged by GPT-4o-mini.
@@
GCG Zou et al. (2023b): Optimizes a discrete adversarial suffix via gradient ascent to maximize compliance likelihood with harmful prompts. GCG does not constrain semantic drift and operates purely at the token level.
@@
Input Embed Schwinn et al. (2024): Performs gradient-based optimization in the continuous embedding space to induce harmful outputs, bypassing token-level defenses.
SCAV Xu et al. (2024): Optimizes for the log-likelihood of harmful completions while explicitly preserving semantic similarity to the original prompt. It includes two variants: SCAV-Embedding, which operates in embedding space, and SCAV-Prompt, which maps optimized embeddings back to discrete tokens.
Weight Orthogonalization Arditi et al. (2024a): A model-level jailbreak method that perturbs the LLM’s feedforward layers by injecting rank-one updates. Specifically, it computes a direction in parameter space orthogonal to existing weights in MLP layers and applies small-scale additive changes. This manipulation increases adversarial compliance by aligning the model’s internal representations with harmful intent.
We use 320 HarmBench samples for GCG, while Prefilling and Input Embed are tested on 240 harmful samples following their original implementation. We run GCG optimization for 500 iterations, with adversarial string initiated as “! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! ! !” (default settings of HarmBench), and searching 512 adversarial candidates at each step. We tested Weight Orthogonalization and SCAV-Embedding on both HarmBench and WidlguardTest samples, and evaluated SCAV-Prompt on AdvBench Chen et al. (2022) samples following the original implementation. The evaluations (except TrustLLM and SCAV-Prompt) use the open-source classifier from HarmBench Mazeika et al. (2024) to determine attack success rates, compliance, and general quality of the responses. TurstLLM-Jailbreak is evaluated with GPT-4.